# ERDŐS–PÓSA PROPERTY OF CYCLES THAT ARE FAR APART

### v1

@@ 3 The Proof
We may assume that $C$ contains no vertex in $\widehat{Y}$ since, otherwise, there is nothing to prove. Since $F_{0}^{-}$ is an induced forest in $G$, $C$ must contain some vertex not in $F_{0}^{-}$. Therefore $C$ contains a vertex in $B_{G}(V(C_{i}),r-d)$ for some $i\in\{1,\ldots,p\}$. By construction, $G[B_{G}(V(C_{i}),r-d)]-Y_{1}$ is unicyclic and, if it contains any cycle then it contains $C_{i}$. Therefore $G[B_{G}(V(C_{i}),r-d)]-(Y_{1}\cup y_{i})$ is a forest. Since $C$ contains no vertex in $Y_{1}\cup\{y_{i}\}\subseteq\widehat{Y}$, $C$ must also contains a vertex not in $B_{G}(V(C_{i}),r-d)$.
Therefore, $C$ contains an edge $v_{0}v_{1}$ with $v_{0}\in B_{G}(V(C_{i}),r-d)$ and $v_{1}\not\in B_{G}(V(C_{i}),r-d)$. If $v_{1}\in B_{G}(V(C_{j}),r)$ for some $j\in[p]\setminus\{i\}$ then $v_{0}v_{1}$ contains an endpoint in $B_{G}(V(C_{i}),r)$ and an endpoint in $B_{G}(V(C_{j}),r)$ in which case at least one of $v_{0}$ or $v_{1}$ is in $Y_{1}$. Since $Y_{1}\subseteq\widehat{Y}$ and $C$ contains no vertex in $\widehat{Y}$, this is a contradiction. Therefore, $v_{1}\not\in B_{G}(V(C_{j}),r-d)$ for all $j\in[p]$. Thus, $v_{1}\in V(F^{-}_{0})$, so $v_{0}v_{1}$ is an $i$-exit edge.
Consider the maximal path $v_{1},\ldots,v_{h-1}$ in $C$ that contains $v_{1}$ and that is contained in $F_{0}^{-}$. Let $v_{h}$ be the neighbour of $v_{h-1}$ in $C-\{v_{h-2}\}$. (Possibly $v_{h}=v_{0}$.) Since $v_{h}\notin\widehat{Y}$ and $v_{h}$ is not a vertex of $F_{0}^{-}$, $v_{h}\in B_{G}(V(C_{j}),r-d)$ for some $j\in\{1,\ldots,p\}$. Then $v_{h-1}v_{h}$ is a $j$-exit edge and $v_{0},\ldots,v_{h}$ is the extended forest leg of the good tuple $t:=(v_{0}v_{1},i,v_{h-1}v_{h},j)$. If $t$ is admissible, then there is nothing more to prove. If $t$ is not admissible, then $B_{G}(V(W_{t}),d)$ contains a vertex of $\widehat{Y}$, say $m$. Let $P_{0}=v_{1},\ldots,v_{h-1}$ be the forest leg of $W_{t}$, let $P_{1}$ be the first leg of $W_{t}$ and let $P_{2}$ be the second leg of $W_{t}$. If $m\in B_{G}(V(P_{1}),d)$ then there is $u\in V(P_{1})$ such that $\operatorname{dist}_{G}(m,u)\leq d$ and so $\operatorname{dist}_{G}(m,v_{0})\leq\operatorname{dist}_{G}(m,u)+\operatorname{dist}_{G}(u,v_{0})\leq d+(r-d)$. If $m\in B_{G}(V(P_{2}),d)$ then $\operatorname{dist}_{G}(m,v_{h})\leq r$. Similarly, if $m\in B_{G}(V(P_{0}),d)$ then $\operatorname{dist}_{G}(m,V(P_{0}))\leq d$. In each case $B_{G}(m,r)$ contains a vertex in $V(C)\supseteq\{v_{0},\ldots,v_{h}\}$. ∎
@@ Acknowledgements
In this appendix, we present a complete proof Theorem 3 and show that, for any fixed $c$, the bounding function $\ell^{\star}(k,c)$ is polynomial in $k$. All of the ideas from this proof come from Gyárfás and Lehel [4], who proved it when the host forest $F$ is a path and sketched changes needed for the general case.

### v2

@@ 3 The Proof
We may assume that $C$ contains no vertex in $\widehat{Y}$ since, otherwise, there is nothing to prove. Since $F_{0}^{-}$ is an induced forest in $G$, $C$ must contain some vertex not in $F_{0}^{-}$. Therefore $C$ contains a vertex in $B_{G}(V(C_{i}),r-d)$ for some $i\in\{1,\ldots,p\}$. By construction, $G[B_{G}(V(C_{i}),r-d)]-Y_{1}$ is unicyclic and, if it contains any cycle then it contains $C_{i}$. Therefore $G[B_{G}(V(C_{i}),r-d)]-(Y_{1}\cup\{y_{i}\})$ is a forest. Since $C$ contains no vertex in $Y_{1}\cup\{y_{i}\}\subseteq\widehat{Y}$, $C$ must also contain a vertex not in $B_{G}(V(C_{i}),r-d)$.
Therefore, $C$ contains an edge $v_{0}v_{1}$ with $v_{0}\in B_{G}(V(C_{i}),r-d)$ and $v_{1}\not\in B_{G}(V(C_{i}),r-d)$. Since $d\geq 1$, $v_{1}\in B_{G}(V(C_{i}),r)$. Now, if $v_{1}\in B_{G}(V(C_{j}),r)$ for some $j\in[p]\setminus\{i\}$, then $v_{1}\in B_{G}(V(C_{j}),r)\cap B_{G}(V(C_{i}),r)\subseteq Y_{1}\subseteq\widehat{Y}$, a contradiction. Therefore, $v_{1}\in V(F_{0}^{-})$, so $v_{0}v_{1}$ is an $i$-exit edge.
Consider the maximal path $v_{1},\ldots,v_{h-1}$ in $C$ that contains $v_{1}$ and that is contained in $F_{0}^{-}$. Let $v_{h}$ be the neighbour of $v_{h-1}$ in $C-\{v_{h-2}\}$. (Possibly $v_{h}=v_{0}$.) Since $v_{h}\notin\widehat{Y}$ and $v_{h}$ is not a vertex of $F_{0}^{-}$, $v_{h}\in B_{G}(V(C_{j}),r-d)$ for some $j\in\{1,\ldots,p\}$. Then $v_{h-1}v_{h}$ is a $j$-exit edge and $v_{0},\ldots,v_{h}$ is the extended forest leg of the good tuple $t:=(v_{0}v_{1},i,v_{h-1}v_{h},j)$. If $t$ is admissible, then there is nothing more to prove. If $t$ is not admissible, then $B_{G}(V(W_{t}),d)$ contains a vertex of $\widehat{Y}$, say $m$. Let $P_{0}=v_{1},\ldots,v_{h-1}$ be the forest leg of $W_{t}$, let $P_{1}$ be the first leg of $W_{t}$ and let $P_{2}$ be the second leg of $W_{t}$. If $m\in B_{G}(V(P_{1}),d)$ then there is $u\in V(P_{1})$ such that $\operatorname{dist}_{G}(m,u)\leq d$ and so $\operatorname{dist}_{G}(m,v_{0})\leq\operatorname{dist}_{G}(m,u)+\operatorname{dist}_{G}(u,v_{0})\leq d+(r-d)$. If $m\in B_{G}(V(P_{2}),d)$ then $\operatorname{dist}_{G}(m,v_{h})\leq r$. Similarly, if $m\in B_{G}(V(P_{0}),d)$ then $\operatorname{dist}_{G}(m,V(P_{0}))\leq d$. In each case $B_{G}(m,r)$ contains a vertex in $V(C)\supseteq\{v_{0},\ldots,v_{h}\}$, a contradiction. ∎
@@ Acknowledgements
In this appendix, we present a complete proof of Theorem 3 and show that, for any fixed $c$, the bounding function $\ell^{\star}(k,c)$ is polynomial in $k$. All of the ideas from this proof come from Gyárfás and Lehel [4], who proved it when the host forest $F$ is a path and sketched changes needed for the general case.